# What About Emotions? Guiding Fine-Grained Emotion Extraction from Mobile App Reviews

### v1

@@ Abstract
Opinion mining plays a vital role in analysing user feedback and extracting insights from textual data. While most research focuses on sentiment polarity (e.g., positive, negative, neutral), fine-grained emotion classification in app reviews remains underexplored. This paper addresses this gap by identifying and addressing the challenges and limitations in fine-grained emotion analysis in the context of app reviews. Our study adapts Plutchik’s emotion taxonomy to app reviews by developing a structured annotation framework and dataset. Through an iterative human annotation process, we define clear annotation guidelines and document key challenges in emotion classification. Additionally, we evaluate the feasibility of automating emotion annotation using large language models, assessing their cost-effectiveness and agreement with human-labelled data. Our findings reveal that while large language models significantly reduce manual effort and maintain substantial agreement with human annotators, full automation remains challenging due to the complexity of emotional interpretation. This work contributes to opinion mining by providing structured guidelines, an annotated dataset, and insights for developing automated pipelines to capture the complexity of emotions in app reviews.
@@ I Introduction
[R1] Useful app which follows Material Design
@@
[R6] Very invasive, […] was forcing to update from a third party store and it want to access everything on your phone including you sim card data
@@
This limitation in expressiveness motivates our research. We propose to use emotion labels to better capture the nuanced emotional states conveyed in user feedback, allowing more informed and targeted decision-making. Fine-grained emotion analysis has been explored in opinion mining tasks in various review types, such as products [16], movies [17] and books [18]. However, emotion analysis research in app reviews remains scarce, particularly in the context of mobile app reviews. This entails several limitations in the state of the art: lack of guidelines for human annotators, ignorance on the challenges they face during the annotation process, and lack of public datasets for multiclass emotion extraction, among others. Moreover, the opportunities large language models (LLMs) offer to perform this task remain also unknown.
@@
    A cost-efficiency and agreement analysis of LLM-based annotations with respect to human agreement.
All datasets and source code are openly shared (See Data Availability Statement at the end of the paper).
@@ II-A Emotion Analysis
Recent advances in natural language processing have sparked growing research interest in opinion mining to support software engineering processes [19]. Moreover, emotional aspects have been extensively explored in the context of RE tasks [20], including elicitation [21, 22, 23], specification [24] and validation [25]. Specifically, analysing user feedback from software-related reviews has also emerged as a key approach to integrating emotional aspects into RE [26, 17, 18]. These studies vary in scope and purpose, ranging from analysing app usage experience [27] to classifying emotional states from user feedback on social media [28].
@@ III-A Design
Figure 1: Research method.
The goal of this research is to identify and address the challenges and limitations of fine-grained emotion analysis in mobile app reviews. Our study focuses on the adaptation of an emotion taxonomy, the analysis of the complexities of human annotation, and the reliability of LLMs for automated annotation. To achieve this goal, we issue the following research questions (RQ):
@@ III-C Iterative Human Annotation
Human annotation takes two key artifacts as input: the emotion taxonomy (derived from RQ1) and the dataset of reviews for annotation. We leveraged a dataset of user reviews from a multi-domain catalogue of popular mobile apps [38], spanning 10 Google Play categories. Additionally, it provides annotations for 198 distinct app features (e.g., instant messaging, video sharing, note-taking, GPS navigation), enabling future research in feature-based emotion analysis. We randomly selected up to 10 sentences from reviews mentioning each feature [^1] Some features in the dataset had fewer than 10 reviews mentioning them., resulting in a total of 1,412 reviews. These were shuffled and divided into 15 subsets, each used in a sequential annotation iteration, labelled from $iteration_{0}$ to $iteration_{14}$. Annotations were performed by five human annotators ($Ann_{1}$ to $Ann_{5}$), all co-authors of this paper, as follows:
@@
  • Step 8 – Dataset annotation. Once the guidelines were stabilized, we iterated over the remaining 12 subsets of reviews ($iteration_{3}\rightarrow iteration_{14}$) to create the final dataset. Three annotators ($Ann_{1},Ann_{2},Ann_{3}$) were designated as main annotators and two ($Ann_{4},Ann_{5}$) as secondary annotators. Each iteration included two main annotators and one secondary annotator, rotating across all possible pairings to systematically identify and address inconsistencies. The 12 iterations were grouped into four batches (2, 4, 4, and 2 iterations, respectively). Each annotator recorded the time spent reading the guidelines and completing each iteration. Each emotion was assessed independently, meaning that a given sentence can be annotated with multiple emotions. This possibility was used with caution, i.e. only when the sentence contained sub-sentences expressing different emotions. The final annotation retained those emotions agreed upon by at least two of the three annotators involved in the iteration.
@@ III-E LLM-based Annotation
After human annotation, we evaluated the performance of LLM-based agents instructed with the generated annotation guidelines. This process was designed as follows:
@@ IV-C Annotation Challenges (RQ3)
Figure 4 summarizes the evolution of the average Cohen’s Kappa agreement for each annotator across iterations following the process described in Sections III-C and III-D. After the initial version of the guidelines produced by $Ann_{1}$ and $Ann_{2}$ in $iteration_{1}$, the team inter-rater agreement grew progressively from moderate at the start of the process (average Cohen’s Kappa of $0.52$ in $iteration_{2}$) to substantial at the end (average Cohen’s Kappa of $0.69$ for all iterations involving the dataset annotation). Notably, all pairwise agreements from $iteration_{3}$ onward remained above the substantial agreement threshold (dotted red line). Fluctuations in early iterations suggest ongoing adjustments into the guidelines, while later stability reflects the effectiveness of iterative refinement. The final iteration showed higher agreement but involved only 12 reviews, making it less representative. Complementarily, we also analysed average Cohen’s Kappa scores for individual emotion labels. Results show that Joy, Anticipation, and Rejectlabels achieved the highest agreement, while Surprise, Anger, and Disgust exhibited the lowest consistency. Expanded label-specific agreement is available in the replication package.
@@ IV-D1 Agreement Analysis
Figure 6 reports the inter-rater pairwise Cohen’s Kappa agreement between all annotators, including five humans [^6] Secondary annotators $Ann_{4}$ and $Ann_{5}$ never participated in the same iteration of the final dataset annotation ($iteration_{3}$ to $iteration_{14}$). ($Ann_{i}$) and LLMs ($GPT_{A}$, $Mistral_{A}$, $Gemini_{A}$). It also includes agreement between each annotator and the overall human ($Ann_{A}$) and LLM-based ($LLM_{A}$) agreements. Agreement between different runs from the same LLM is excluded from this analysis, as they exhibited near-perfect consistency (pairwise Cohen’s Kappa agreement was always $\geq 0.87$). On average, humans show similar agreement among themselves ($0.69$) as LLMs ($0.68$). However, LLM annotations ($GPT_{A}$, $Mistral_{A}$, $Gemini_{A}$) exhibit moderate agreement with human annotations ($0.56-0.62$), indicating some deviation from human judgment. Among individual models, Gemini 2.0 Flash ($Gemini_{A}$) aligns more closely with human agreement ($Ann_{A}$), although LLM agreement ($LLM_{A}$) is slightly higher. This suggests that combining multiple LLM instances helps mitigate biases, similar to how human annotators vary in judgment. GPT-4o ($GPT_{A}$) shows the highest consistency with overall LLM agreement ($LLM_{A}$), while Mistral and Gemini introduce more variability across different runs.
TABLE II: LLM correctness vs. human ($Ann_{A}$)
@@
TABLE III: LLM agreement ($LLM_{A}$) correctness per annotation label
@@ IV-D2 Cost-efficiency Analysis
Table IV compares the cost-efficiency of human and LLM-based annotators [^7] Extended results per annotator are included in the replication package.. For better comparability, the table reports average time and cost per annotator. For human annotators ($Ann_{i}$), we compute the average across the three annotators participating in each iteration. For LLM annotators ($GPT_{i}$, $Mistral_{i}$, $Gemini_{i}$), we take the average across the three annotation runs. On average, a human annotator takes more than 7×7$\times$7 × longer to annotate the whole dataset ($\approx 7$ hours) than the slowest LLM-based annotator [^8] LLM annotators integrate annotation guidelines as system prompts during each iteration; therefore, guideline processing time is not reported., Gemini ($\approx 1$ hour). Cost disparity is even more pronounced – on average, one human annotator is 233×233$\times$233 × more expensive [^9] Based on the average hourly cost of the five annotators. than the most costly LLM annotator, GPT-4o. These findings highlight the scalability of LLM-based annotation, offering a cost-effective alternative to manual annotation, particularly for iterative and large-scale annotation tasks.
TABLE IV: Cost-efficiency analysis (average per annotator)
@@
Time Anni 11.4’ 38.4’ 427.0’
GPTi - 2.9’ 31.9’
Mistrali - 2.2’ 24.7’
Geminii - 5.2’ 57.9’
@@ V-A3 What challenges arise when annotating app reviews with emotional labels? (RQ3)
V-B Threats to Validity
@@
Concerning internal validity [62], annotators relied on the full review to resolve linguistic ambiguities, which may have introduced unintended dependencies on contextual information. While this approach improved consistency, it also introduced the risk that certain emotional cues may have been interpreted differently when considering broader review contexts. Additionally, personal bias during human annotation posed a threat, particularly when annotators encountered ambiguous or overlapping emotional expressions. While involving five annotators and the generation, discussion, and refinement of the guidelines reduced this risk, differences in annotator interpretations may still have led to slight variations in the annotated dataset. Furthermore, annotation challenges such as distinguishing subtle variations in intensity and resolving conflicting emotions (e.g., Joy vs. Trust, Sadness vs. Disgust) may have affected annotation reliability. To address this, we specifically addressed these challenges through dedicated discussions and mitigation actions to improve the guidelines. Furthermore, we report these challenges to assist further research in considering the limitations of our dataset properly.
@@
For construct validity, the selection of Plutchik’s taxonomy might not fully capture all possible emotional states present in app reviews. We mitigated this risk by conducting a thorough literature review before selecting the taxonomy. In addition, as our research design remains agnostic to a specific taxonomy, our replication package enables adaptation to alternative emotion taxonomies and annotation schemes. Another limitation stemming from the annotation process (both human and LLM) is the assessment of emotional intensity, where subtle variations in arousal remain difficult to measure consistently.
@@
Finally, concerning external validity, our dataset may not generalize to all app review domains. Although reviews span diverse categories and features, further validation is needed to assess applicability across other domains, languages, and cultural contexts. Linguistic differences and user demographics may affect how emotions are conveyed in app reviews, suggesting the need for cross-domain evaluation. Regarding LLM selection, our results may not fully generalize to other architectures or newer models. Exploring alternatives like DeepSeek [^10] https://www.deepseek.com/ (with restricted API access during this research) or OpenAI reasoning models (e.g., o1), now available via API, could yield different outcomes, potentially improving agreement or introducing new annotation biases.

### v2

@@ Abstract
Opinion mining plays a vital role in analysing user feedback and extracting insights from textual data. While most research focuses on sentiment polarity (e.g., positive, negative, neutral), fine-grained emotion classification in app reviews remains underexplored. Fine-grained emotion classification is thus needed to better understand users’ affective responses and support downstream tasks such as feature-emotion analysis, user-oriented release planning, and issue triaging. This paper addresses this gap by identifying and addressing the challenges and limitations in fine-grained emotion analysis in the context of app reviews. Our study adapts Plutchik’s emotion taxonomy to app reviews by developing a structured annotation framework and dataset. Through an iterative human annotation process, we define clear annotation guidelines and document key challenges in emotion classification. Additionally, we evaluate the feasibility of automating emotion annotation using large language models, assessing their cost-effectiveness and agreement with human-labelled data. Our findings reveal that while large language models significantly reduce manual effort and maintain substantial agreement with human annotators, full automation remains challenging due to the complexity of emotional interpretation. This work contributes to opinion mining in requirements engineering by providing structured guidelines, an annotated dataset, and insights for developing automated pipelines to capture the complexity of emotions in app reviews.
@@ I Introduction
[R1] Useful app which follows Material Design.
@@
[R6] Very invasive, […] was forcing to update from a third party store and it wanted to access everything on your phone including your SIM card data.
@@
This limitation in expressiveness motivates our research. We propose to use emotion labels to better capture the nuanced emotional states conveyed in user feedback, allowing more informed and targeted decision-making. This enables use cases such as fine-grained feature-oriented emotion analysis to better prioritize or refine specific app functionalities, and market analysis to categorize user feedback more effectively by emotional trends in terms of user satisfaction and adoption.
Fine-grained emotion analysis has been explored in opinion mining tasks in various review types, such as products [16], movies [17] and books [18]. However, emotion analysis research in app reviews remains scarce, particularly in the context of mobile app reviews. This entails several limitations in the state of the art: lack of guidelines for human annotators, ignorance on the challenges they face during the annotation process, and lack of public datasets for multiclass emotion extraction, among others. Moreover, the opportunities large language models (LLMs) offer to perform this task remain unknown.
@@
    Insights into the cost-effectivenes, agreement and correctness of LLM-based annotations with respect to humans.
A replication package including datasets and source code is openly shared (See Data Availability Statement).
@@ II-A Emotion Analysis
Recent advances in natural language processing have sparked growing research interest in opinion mining to support software engineering processes [19]. Moreover, emotional aspects have been extensively explored in the context of RE tasks [20], including elicitation [21, 22, 23], specification [24] and validation [25]. Specifically, analysing user feedback from software-related reviews has also emerged as a key approach to integrating emotional aspects into RE [26, 17, 18]. These studies vary in scope and purpose, ranging from analysing app usage experience [27] to classifying emotional states from user feedback on social media platforms [28, 29].
@@ III-A Design
The goal of this research is to identify and address the challenges and limitations of fine-grained emotion analysis in mobile app reviews. Our study focuses on the adaptation of an emotion taxonomy, the analysis of the complexities of human annotation, and the reliability of LLMs for automated annotation. The following use cases illustrate how this fine-grained emotion analysis can support RE-related feedback gathering tasks by uncovering insights beyond traditional polarity-based analysis.
  1. UC1.
    Categorizing feature-related feedback. Identify distinct emotional responses tied to specific features, such as whether users struggle with usability, expect improvements, express confusion, or show appreciation.
  2. UC2.
    Emotion-based user segmentation for release planning. Segment user feedback by emotion types to understand how different user groups emotionally respond to recent updates or unmet expectations. This supports planning future releases that align better with user sentiment dynamics across versions or personas.
  3. UC3.
    Prioritizing critical issues based on emotional risk. Analyse high arousal reviews (e.g., anger over privacy violations, fear about data loss, or betrayal due to broken trust) to prioritize and address issues with higher potential for reputational or legal consequences.
To achieve this goal and support the outlined use cases, we issue the following research questions (RQ):
@@ III-C Iterative Human Annotation
Human annotation takes two key artifacts as input: the emotion taxonomy (derived from RQ1) and the dataset of reviews for annotation. We leveraged a dataset of user reviews from a multi-domain catalogue of popular mobile apps [39], spanning 10 Google Play categories. Additionally, it provides annotations for 198 distinct app features (e.g., instant messaging, video sharing, note-taking, GPS navigation), enabling future research in feature-based emotion analysis. We randomly selected up to 10 sentences from reviews mentioning each feature [^1] Some features in the dataset had fewer than 10 reviews mentioning them., resulting in a total of 1,412 reviews. These were shuffled and divided into 15 subsets (100 reviews each, 12 for the latter), labelled from $iteration_{0}$ to $iteration_{14}$. Annotations were performed by five human annotators ($Ann_{1}$ to $Ann_{5}$), all co-authors of this paper, as follows:
@@
  • Step 8 – Dataset annotation. Once the guidelines were stabilized, we iterated over the remaining 12 subsets of reviews ($iteration_{3}\rightarrow iteration_{14}$) to create the final dataset. Three annotators ($Ann_{1},Ann_{2},Ann_{3}$) were designated as main annotators and two ($Ann_{4},Ann_{5}$) as secondary annotators. Each iteration included two main annotators and one secondary annotator, rotating across all possible pairings to systematically identify and address inconsistencies. The 12 iterations were grouped into four batches (2, 4, 4, and 2 iterations, respectively). Each annotator recorded the time spent reading the guidelines and completing each iteration. Each emotion was assessed independently, meaning that a given sentence can be annotated with multiple emotions. This possibility was used with caution, i.e. only when the sentence contained sub-sentences expressing different emotions. The final annotation retained emotion labels that at least two of the three annotators involved in the iteration agreed upon.
@@ III-E LLM-based Annotation
After annotation, we evaluated LLM agents using the generated guidelines as follows:
@@ IV-C Annotation Challenges (RQ3)
Figure 4 summarizes the evolution of the average Cohen’s Kappa agreement for each annotator across iterations following the process described in Sections III-C and III-D.
@@
After the initial version of the guidelines produced by $Ann_{1}$ and $Ann_{2}$ in $iteration_{1}$, the team inter-rater agreement grew progressively from moderate at the start of the process (average Cohen’s Kappa of $0.52$ in $iteration_{2}$) to substantial at the end (average Cohen’s Kappa of $0.69$ for all iterations involving the dataset annotation). Notably, all pairwise agreements from $iteration_{3}$ onward remained above the substantial agreement threshold (dotted red line). Fluctuations in early iterations suggest ongoing adjustments into the guidelines, while later stability reflects the effectiveness of iterative refinement. The final iteration showed higher agreement but involved only 12 reviews, making it less representative. Complementarily, we also analysed average Cohen’s Kappa scores for individual emotion labels. Results show that Joy, Anticipation, and Rejectlabels achieved the highest agreement, while Surprise, Anger, and Disgust exhibited the lowest consistency. Expanded label-specific agreement is available in the replication package.
@@ IV-D1 Agreement Analysis
Figure 6 reports the inter-rater pairwise Cohen’s Kappa agreement between all annotators, including five humans [^6] Secondary annotators $Ann_{4}$ and $Ann_{5}$ never participated in the same iteration of the final dataset annotation ($iteration_{3}$ to $iteration_{14}$). ($Ann_{i}$) and LLMs ($GPT_{A}$, $Mistral_{A}$, $Gemini_{A}$). It also includes agreement between each annotator and the overall human ($Ann_{A}$) and LLM-based ($LLM_{A}$) agreements. Agreement between different runs from the same LLM is excluded from this analysis, as they exhibited near-perfect consistency (pairwise Cohen’s Kappa agreement was always $\geq 0.87$). On average, pairs of humans show an agreement ($0.69$) that is similar to the agreement shown by pairs of LLMs ($0.68$). However, LLM annotations ($GPT_{A}$, $Mistral_{A}$, $Gemini_{A}$) exhibit moderate agreement with human annotations ($0.56-0.62$), indicating some deviation from human judgment. Among individual models, Gemini 2.0 Flash ($Gemini_{A}$) aligns more closely with human agreement ($Ann_{A}$), although LLM agreement ($LLM_{A}$) is slightly higher. This suggests that combining multiple LLM instances helps mitigate biases, similar to how human annotators vary in judgment. GPT-4o ($GPT_{A}$) shows the highest consistency with overall LLM agreement ($LLM_{A}$), while Mistral and Gemini introduce more variability across different runs.
TABLE II: LLM correctness vs. human agreement ($Ann_{A}$)
@@
TABLE III: Human individual correctness vs. human agreement ($Ann_{A}$)
Annotator Precision Recall F1
Ann1 86.96% 88.76% 87.86%
Ann2 88.66% 88.89% 88.78%
Ann3 86.14% 84.84% 85.49%
Ann4 85.24% 83.08% 84.16%
Ann5 81.97% 77.73% 79.85%
Anni 85.79% 84.66% 85.23%
Complementarily, and using the same criteria, Table III shows that human annotators are highly consistent with one another, with F1 scores ranging narrowly from 79.85% to 88.78% and an overall average of 85.23%. Despite having different roles in the annotation process (e.g., primary vs. secondary), all annotators achieve comparably high precision and recall, indicating strong alignment with the agreed-upon annotation guidelines. This consistency reinforces the reliability of the human-created ground truth and highlights the effectiveness of the annotation guidelines and calibration process in minimizing individual biases.
@@
TABLE IV: LLM agreement ($LLM_{A}$) correctness per label
@@
TABLE V: Average human individual ($Ann_{i}$) correctness per label
Label Precision Recall F1
Joy 90.99% 91.37% 91.18%
Sadness 89.36% 82.67% 86.02%
Anticipation 88.71% 88.72% 88.71%
Trust 85.23% 81.37% 83.30%
Surprise 64.07% 74.56% 69.31%
Disgust 68.98% 72.62% 70.80%
Anger 69.07% 77.21% 73.14%
Fear 74.15% 78.50% 76.32%
Neutral 78.41% 76.06% 77.24%
Reject 82.00% 89.46% 85.73%
Table V shows the average correctness of individual human annotators per emotion label. While humans also find emotions like Surprise, Disgust, Anger, and Fear more challenging, their performance remains stable and never drops to the critical levels seen in LLM outputs. In contrast, LLMs significantly underperform on these emotions, with pronounced precision-recall imbalances and frequent misclassifications. This suggests that LLMs struggle with subtle or ambiguous emotions that humans, despite some variation, handle more reliably.
@@ IV-D2 Cost-efficiency Analysis
Table VI compares the cost-efficiency of human and LLM-based annotators [^7] Extended results per annotator are included in the replication package.. For better comparability, the table reports average time and cost per annotator. For human annotators ($Ann_{i}$), we compute the average across the three annotators participating in each iteration. For LLM annotators ($GPT_{i}$, $Mistral_{i}$, $Gemini_{i}$), we take the average across the three annotation runs. On average, a human annotator takes more than 7×7$\times$7 × longer to annotate the whole dataset ($\approx 7$ hours) than the slowest LLM-based annotator [^8] LLM annotators integrate annotation guidelines as system prompts during each iteration; therefore, guideline processing time is not reported., Gemini ($\approx 1$ hour). Cost disparity is even more pronounced – on average, one human annotator is 233×233$\times$233 × more expensive [^9] Based on the average hourly cost of the five annotators. than the most costly LLM annotator, GPT-4o. These findings highlight the scalability of LLM-based annotation, offering a cost-effective alternative to manual annotation, particularly for iterative and large-scale annotation tasks.
TABLE VI: Cost-efficiency analysis (average per annotator)
@@
Time Anni 11.4 38.4 427.0
GPTi - 2.9 31.9
Mistrali - 2.2 24.7
Geminii - 5.2 57.9
@@ V-A3 What challenges arise when annotating app reviews with emotional labels? (RQ3)
V-B Research Goal and Implications
Our study showcases how fine-grained emotion analysis in mobile app reviews faces three core limitations: (1) the lack of domain-adapted taxonomies, (2) the ambiguity in human interpretation of emotion labels, and (3) the uncertain reliability of automated methods. Our study directly addresses these limitations by operationalizing a widely accepted taxonomy (i.e., Plutchik’s) for the mobile app domain, developing annotation guidelines that clarify emotion boundaries with domain-specific examples, and empirically demonstrating where LLMs align or diverge from human annotation. These contributions not only make emotion classification more feasible and interpretable but also expose the practical constraints and design trade-offs in scaling annotation processes—offering grounded insight for future tool development and dataset construction.
As a result, our study unlocks new opportunities in traditional RE tasks within the context of opinion mining from app stores. Our contributions (i.e., annotation guidelines, dataset, automated annotation pipelines) serve as groundwork for building (semi-)automatic pipelines for more precise categorization of user feedback, enabling a diversity of use cases that can be explored by leveraging the artifacts resulting from this research—such as feature-emotion pair sentiment analysis (UC1), emotionally-oriented user segmentation for release planning (UC2), and fine-grained issue prioritization (UC3). The analysis of the generated dataset (RQ2) and the challenges identified during its generation (RQ3) demonstrate how emotional labels such as Anticipation, Trust, Surprise, or Fear offer extended knowledge dimensions compared to traditional polarity-based approaches. Furthermore, emotion analysis not only conveys the insights captured by polarity (i.e., positive, negative) and review type (i.e., bug report, feature request, praise [3]) classification pipelines, but also expands on additional emotional dimensions (e.g., Surprise, Fear, Trust) that might otherwise be neglected.
V-C Threats to Validity
@@
Concerning threats to validity [63], annotators relied on the full review to resolve linguistic ambiguities, which may have introduced unintended dependencies on contextual information. While this approach improved consistency, it also introduced the risk that certain emotional cues may have been interpreted differently when considering broader review contexts. Additionally, personal bias during human annotation posed a threat, particularly when annotators encountered ambiguous or overlapping emotional expressions. While involving five annotators and the generation, discussion, and refinement of the guidelines reduced this risk, differences in annotator interpretations may still have led to slight variations in the annotated dataset. Furthermore, annotation challenges such as distinguishing subtle variations in intensity and resolving conflicting emotions (e.g., Joy vs. Trust, Sadness vs. Disgust) may have affected annotation reliability. To address this, we specifically addressed these challenges through dedicated discussions and mitigation actions to improve the guidelines. Furthermore, we report these challenges to assist further research in considering the limitations of our dataset properly.
@@
The selection of Plutchik’s taxonomy might not fully capture all possible emotional states present in app reviews. We mitigated this risk by conducting a thorough literature review before selecting the taxonomy. In addition, as our research design remains agnostic to a specific taxonomy, our replication package enables adaptation to alternative emotion taxonomies and annotation schemes. Another limitation from the annotation process is the assessment of emotional intensity, as subtle arousal variations remain difficult to measure consistently.
@@
Finally, our dataset may not generalize to all app review domains. Although reviews span diverse categories and features, further validation is needed to assess applicability across other domains, languages, and cultural contexts. Linguistic differences and user demographics may affect how emotions are conveyed in app reviews, suggesting the need for cross-domain evaluation. Regarding LLM selection, our results may not fully generalize to other architectures or newer models. Exploring alternatives like DeepSeek (with restricted API access during this research) or OpenAI reasoning models (e.g., o1), now available via API, could yield different outcomes, potentially improving agreement or introducing new annotation biases.